OpenStudy (anonymous):

What is the Y-Intercept of the equation below? 4y = 3x

OpenStudy (alexandervonhumboldt2):

convert to slope intercept form first

OpenStudy (alexandervonhumboldt2):

in slope intercept form yax+b b is the y-inter

OpenStudy (anonymous):

Convert to slope-intercept form. y = mx + b 4y = 3x Divide 4 to both sides: y = 3/4x y = mx + b Where 'b' is the y-intercept, so what's the y-intercept here?

OpenStudy (nincompoop):

in the slope-inperecept form \(y=mx+\color{red}{b},\) the b is the y-intercept.

OpenStudy (anonymous):

y = 3/4x Do you see anything after '3/4x'? @tmowry228

OpenStudy (anonymous):

no

OpenStudy (nincompoop):

remember that \(y=mx+\color{red}{0} \) is also the same as saying that the y-intercept is zero, which we normally omit such as in the case of \(y = 3x+\color{red}{0} \rightarrow y=3x\)

OpenStudy (alexandervonhumboldt2):

y=ax+b int his case b=0

OpenStudy (holly00d1248):

then what would it be nothing eequals what

OpenStudy (nincompoop):

that is the rough way of looking at it, but it is important to note that the y-intercept can actually be analytically solved by making the value of \(x=0 \), which will make sense if you have an idea how to graph linear equations

OpenStudy (anonymous):

Yes, then the y-intercept is 0.

OpenStudy (anonymous):

\(\bbox [10pt, lime, border:5pt solid black]{\Huge\cal\color{red}\diamondsuit\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\diamondsuit}}}\)